Try It 1.2.2 Find a conversion factor between square centimeters and square meters. Write it in three forms. Complete the conversion factor below. centimeters 1 square meter Enter your answer in the answer box and then click Check Answer. Clear All Check Answer remaining
Answer #1

1 meter = 100 cm

1m^2 = 100^2 cm^2

1m^2 = 10000 cm^2

1m^2 = 10^4 cm^2

2 Im x Im 2. 2 2.